OpenStudy (anonymous):

quick question we have \(ab=cd\) if \(\gcd (a,b)=1\) , \(\gcd (c,d)=1\) ...\((a

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

im confused...help

OpenStudy (anonymous):

i have to think for a second but i would say no

OpenStudy (anonymous):

this just says \(a\) and \(b\) are made up of distinct primes, i.e. they have no primes in common

OpenStudy (anonymous):

similarly so are \(c\) and \(d\)

OpenStudy (anonymous):

so for example \(7\times 15=35\times 3\)

OpenStudy (anonymous):

i see that i ignored your inequality, but no matter \[7\times 15=3\times 35\]

OpenStudy (anonymous):

lol....:D....thank u...what a question...!!!!

OpenStudy (anonymous):

yw. morning soft ball